OpenStudy (katiewayne1975):

What is the equation of the line x–3y=18 in slope-intercept form? Drag and drop the appropriate number, symbol, or variable to each box.

OpenStudy (tgstudios):

welcome to openstudy!

OpenStudy (katiewayne1975):

@triciaal can you help me please

HanAkoSolo (jamierox4ev3r):

first of all, can you tell me what slope-intercept form is?

HanAkoSolo (jamierox4ev3r):

Like without any numbers whatsoever, what does the slope-intercept form look like?

HanAkoSolo (jamierox4ev3r):

@KatieWayne1975

OpenStudy (katiewayne1975):

x-3y=18 this is it

OpenStudy (katiewayne1975):

@ShadowLegendX

HanAkoSolo (jamierox4ev3r):

you want to manipulate the equation so that you are effectively solving for y

OpenStudy (mathmale):

that was a good question. Jamie wants to be sure you know what form your answer must have. It's "slope-intercept form," so must look like y=mx+b. m is the slope and b is the y-intercept. Your x-3y=18 is in Standard Form, or close to it; your job is to convert this to Slope-Intercept form." Hint: just solve x-3y=18 for y. Show your work.

OpenStudy (will.h):

Slope intercept form is. Y = mx + b So.isolate y and you'll get your expression

HanAkoSolo (jamierox4ev3r):

@KatieWayne1975 do you understand what we're wanting you to do? So you are given x-3y=18. We want you to find y, because the slope-intercept form is this: y=mx+b, where m and b are simply numbers

OpenStudy (mathmale):

correct, jamie, but do you realize that part of what you're sharing here simply repeats information given out earlier?

HanAkoSolo (jamierox4ev3r):

I was just trying to connect it all together, hopefully make it clearer.

HanAkoSolo (jamierox4ev3r):

I'm not sure what more I can do, as this problem involves basic variable solving at this point.